When 47.4 g of KClO3  is completely decomposed, what is the theoretical yield, in grams, of O2 ?     


1
Expert's answer
2022-02-19T06:52:56-0500

The equation for the decomposition is;

2KClO3(s) "\to"2KCl(s) + 3O2(g)

Moles of KClO3 = 47.4÷122.5

=0.3869

Moles of oxygen=(0.3869÷2)×3

=0.58035

Yield of oxygen = 0.58035 × 32

= 18.5712g